OpenStudy (anonymous):

use mental math to find each product 1. 5.1 times 8 2. 299 times 3 3.9 times 6 4.6.15 times 4

OpenStudy (kymber):

You're supposed to use mental math!

OpenStudy (anonymous):

1. 5*8 +.1*8=40.8 2. 300*3 - 1*3 = 297 3. 10*6 - 1*6 = 54 4. 6.15*4 = 6*4 +.15*4=24.6

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

2. is 897